OpenStudy (liljj1421):

explain how to write a function rule from the table below. then write a function rule. |x|2|4|6| |y|1|0|-1|

OpenStudy (liljj1421):

which graph represents the liner function y= 1/4*-3?

OpenStudy (anonymous):

its the 2nd graph

OpenStudy (liljj1421):

wich graph represents the linear function y=1/3*x-4

OpenStudy (anonymous):

put x =0 => y=-4

OpenStudy (anonymous):

then put y=0 => x=+12

OpenStudy (anonymous):

c

OpenStudy (liljj1421):

witch graph represents the linear function y=3*x+1

OpenStudy (liljj1421):

sorry these ones go to the question above

OpenStudy (liljj1421):

@HackberryAbby18

OpenStudy (liljj1421):

@Thesmarterone @HackberryAbby18 @grex10

OpenStudy (liljj1421):

I need help on the first and last question.

OpenStudy (liljj1421):

@Omar_Elboredy

OpenStudy (liljj1421):

please

OpenStudy (liljj1421):

@prettypink_mega

OpenStudy (liljj1421):

explain how to write a function rule from the table below. then write a function rule. |x|2|4|6| |y|1|0|-1

OpenStudy (liljj1421):

witch graph represents the linear function y=3*x+1 4 attachments 913045731201394608am7899723053.jpg 913045731201394608am7899723053.jpg 913045731201394608am13492647082.jpg 913045731201394608am13492647082.jpg 913045731201394608am15533028022.jpg 913045731201394608am15533028022.jpg 869973731201315229pm4962020422.jpg 869973731201315229pm4962020422.jpg

OpenStudy (anonymous):

I 'll write step-by-step but stay calm ! :)

OpenStudy (liljj1421):

I just need help with the bottom two

OpenStudy (anonymous):

K, you know equation of line ?

OpenStudy (liljj1421):

no

OpenStudy (anonymous):

K, dun worry , first you need to know that any line has an equation of the form : \[y = mx +c\] ; where m represents the slope of that line and c is the intercept part with the y-axis , K ?